OpenStudy (anonymous):

f(x)= 1/3 x^3+1/2 x ^2 x^2-6x+1 the function is only valid in the domain -5 bigger or qual = x = smaller or equal 5 how do you find the maxima and the minima?

OpenStudy (anonymous):

@wilbert06 Do you know how to find critical points?

OpenStudy (anonymous):

reminder: derivatives

OpenStudy (anonymous):

no clue

OpenStudy (anonymous):

i know how to derrive but i dont understand the question

OpenStudy (anonymous):

This is actually not that hard, I'll help. First, take your derivative. Find f'(x). Type it in here when you've got it, ok?

OpenStudy (anonymous):

x^2 +x -6x^-1+ 0

OpenStudy (anonymous):

f'(x) =x^2 +x -6x^-1+ 0

OpenStudy (anonymous):

@agentx5 there you have it what do i do next?

OpenStudy (anonymous):

Ah so not \[\frac{d}{dx}(x^3/3+(x^2 x^2)/2-6 x+1) = 2 x^3+x^2-6\] You had this as your original then? \[\int\limits (x^2+x-6/x+0) dx = 1/6 x^2 (2 x+3)-6 log(x)+constant\] (integral goes in reverse)

OpenStudy (anonymous):

I don't think I have your original function quite understood. Show me with parenthesis what you meant.

OpenStudy (anonymous):

ok let me try to write it doen better

OpenStudy (anonymous):

The next step will be to evaluate for 0's, because that's when the SLOPE is 0, or it's maxed or minimized out. Make sense? It's got no slope at the tip of a peak or valley.

OpenStudy (anonymous):

\[f(x)= 1/3 x ^{3}+ 1/2^{2}-6x+1\]

OpenStudy (anonymous):

that was the given fucntion

OpenStudy (anonymous):

so step one would be differentiate it?

OpenStudy (anonymous):

\[1/2^2+1=\frac{1}{4}+1=\frac{5}{4}\] So that's just your constant, yeah? And yes, step one is exactly that

OpenStudy (anonymous):

so after its differentiated you solve x, that means you make the equation = 0?

OpenStudy (anonymous):

Check to see if you have this: \[ \frac{d}{dx}(\frac{x^3}{3}-6 x+\frac{5}{4}) = x^2-6 \]

OpenStudy (anonymous):

Then you'll have your f'(x) function. The critical points are f'(0), plug in zero. As you might be able to see because it's a linear function, you're only going to have one point.

OpenStudy (anonymous):

yes its that

OpenStudy (anonymous):

ahhh i see the zero is when the slope is a straight line

OpenStudy (anonymous):

Take the derivative again of x^2-6 and see if it's +, -, or 0. This tells you the concavity. It's just 2, which is a + number so it opens upward there.

OpenStudy (anonymous):

Yeah you'll eventually be able to do this in your head, and in Calc2 or Calc3 you'll really need to in fact.

OpenStudy (anonymous):

Here let me give you the graph and you'll be facepalming :D

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

See if while I'm grabbing this you can find coordinate for the original function for that point when the slope is zero. There's more than one because it's an x cubed function originally

OpenStudy (anonymous):

It's going to cross the x axis in 3 locations, 3 roots

OpenStudy (anonymous):

ok so again, step 1 would be differentiate, step 2 would be solve x for zero and then?

OpenStudy (anonymous):

the original would be the integral?

OpenStudy (anonymous):

derivative and integral are inverse operations just like... addition and subtraction multiplication and division exponent and logarithm And here's your graph to look at:

OpenStudy (anonymous):

Forward & reverse, basically. This is a core concept of ALL mathematics. If you can do it one way, you better be able to reverse it with the right conditions.

OpenStudy (anonymous):

Technically it's derivative and anti-derivative and the anti-derivative with the conditions make it into the operation called integral

OpenStudy (anonymous):

ok aha

OpenStudy (anonymous):

What you're doing to find critical points is finding when slope is zero. And in this case the GLOBAL minimum and maximum are -∞ and ∞

OpenStudy (anonymous):

You were asked to find local I presume, which is those peaks or valleys "bends" when it changes direction.

OpenStudy (anonymous):

ohh ok local and global theres a difference

OpenStudy (anonymous):

Not always. But for this kind of function? yes.

OpenStudy (anonymous):

yeah i guess the local you know between 2 and 4 in think

OpenStudy (anonymous):

they did ask between -5 and 5

OpenStudy (anonymous):

Between 5 and -5 is your domain, what they're restricting you to. Major hint there it's local min's or max's in that section

OpenStudy (anonymous):

Graph the parabola y=x^2-6 and you'll find where it crosses the x-axis (when (f'(x)=0) is EXACTLY where the slope is zero

OpenStudy (anonymous):

abc formula not aply able?

OpenStudy (anonymous):

square root of 6?

OpenStudy (anonymous):

ax^2+bx+c? Um, no that's only for special cases of polynomials

OpenStudy (anonymous):

Yes, it's the squareroot of 6, plus or minus

OpenStudy (anonymous):

The local max is at the minus, the local min is at the plus

OpenStudy (anonymous):

ohhh !

OpenStudy (anonymous):

and wala that was it?

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

Don't forget to specific coordinates though

OpenStudy (anonymous):

Got to answer the original problem, but now you know what x has to be in order for slope to be zero for the original problems and which is which

OpenStudy (anonymous):

so the answer would be ,step 1) differentiate step 2) find y=0 step 3) is the answer?

OpenStudy (anonymous):

y'=0 , there's a diff ;-)

OpenStudy (anonymous):

haha ok

OpenStudy (anonymous):

Not trying to be nitpicky but it's kinda important to note it's the derivative

OpenStudy (anonymous):

yes yes

OpenStudy (anonymous):

(x , y )

OpenStudy (anonymous):

ah sooo the specific coordinates

OpenStudy (anonymous):

i would have to fill in the square root of 6 in th original formula?

OpenStudy (anonymous):

@agentx5 is that so?

OpenStudy (anonymous):

@agentx5 thanks for the help! i see it now atleast i think i understand the method

OpenStudy (anonymous):

@agentx5 goodnight